OpenStudy (anonymous):

Solve the equation. Check the solution. Please show steps. Please help me. :( 1/b+1 + 1/b-1 = 2/b^2-1

OpenStudy (anonymous):

@Shawnee01 will you help me? :(

OpenStudy (gorv):

\[\frac{ 1 }{ b+1 }+\frac{ 1 }{ b-1 }=\frac{ 2 }{b ^{2}-1 }\]

OpenStudy (gorv):

is it ur Q

OpenStudy (gorv):

take L.C.M of left side

OpenStudy (gorv):

\[\frac{ (b+1)+(b-1) }{ (b+1)(b-1) }=\frac{ 2b }{ b^2-1 }\]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

is that it?

OpenStudy (gorv):

left side equals right side that's it

OpenStudy (anonymous):

Thankyou , Gorv

OpenStudy (gorv):

welcme @Emily778